ELECTRONIC TRANSITIONS OF PALLADIUM AND VANADIUM DIMER

\underline{YUE QIAN}, Y.W. NG, and A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

The laser induced fluorescence (LIF) spectrum of palladium dimer (Pd_{2}) in the visible region between 480 and 700 nm has been studied. Five vibrational bands were recorded and analyzed; they are assigned to a ^{3}Π_{g} - X^{3}Σ_{u}^{+} system. The vibrational frequency of the ground X^{3}Σ_{u}^{+} state has been determined to be 211.4 cm^{−1}. This is the first experimental observation of the LIF spectrum of Pd_{2}. In addition, the LIF spectrum of vanadium dimer (V_{2}) has also been studied; several new transition band systems were observed in the wavelength between 480 and 530 nm. The analysis of the spectra recorded for these two molecules will be presented.